# e-Profits: A Business-Aligned Evaluation Metric for Profit-Sensitive Customer Churn Prediction

### v1

@@ 6 Conclusion
Traditional churn prediction models are often assessed using statistical performance metrics such as Accuracy, AUC, and F1-score. While these metrics serve well in quantifying classification performance, they frequently fail to capture the real financial consequences of misclassification, particularly in profit-sensitive applications like customer retention.
In this study, we introduced e-Profits, a business-aligned evaluation metric that integrates customer-specific lifetime value (CLV), intervention costs, and retention probabilities derived from Kaplan-Meier survival analysis. Unlike traditional metrics and even profit-based alternatives such as EMP, e-Profits evaluates models based on individualised cost-revenue dynamics, facilitating better alignment between predictive analytics and financial outcomes. Our empirical evaluation on two benchmark datasets demonstrated that e-Profits significantly reshapes classifier rankings. Models that were suboptimal under Accuracy or AUC often emerged as the most financially effective under e-Profits, particularly when focusing on the top 20% of customers by predicted profitability. This shift underscores the potential for misaligned decision-making if only conventional metrics are used. The metric’s post hoc nature and lightweight implementation make e-Profits practical for real-world deployment. It requires no retraining and can be integrated easily into existing evaluation workflows, making it accessible to business analysts and operational stakeholders alike. Moreover, its understandable structure enhances communication between technical teams and decision-makers.
Future directions for e-Profits include embedding it as an objective function during model training and experimenting with advanced survival models for improved retention prediction. Overall, e-Profits introduces a financially meaningful evaluation paradigm that empowers practitioners to make data-driven decisions rooted in return on investment.
CRediT authorship contribution statement
Awais Manzoor: Conceptualization, Data curation, Methodology and Experimental setup, Software, Formal analysis, Investigation, Validation, Writing – original draft, Writing – review & editing M. Atif Qureshi: Conceptualization, Data curation, Methodology and Experimental setup, Software, Formal analysis, Investigation, Validation, Writing – review & editing, Supervision Etain Kidney: Supervision, Writing – review & editing Luca Longo: Supervision, Writing – review & editing
Declaration of competing interest

### v2

@@ 6 Conclusion
Traditional churn prediction models are commonly evaluated using statistical performance metrics such as Accuracy, AUC, and F1-score. While these metrics are useful for assessing classification performance, they often fail to reflect the financial consequences of misclassification, particularly in profit-sensitive applications such as customer retention.
In this paper, we introduced e-Profits, a business-aligned evaluation metric that integrates customer-specific lifetime value, intervention costs, and retention probabilities derived from Kaplan-Meier survival analysis. Unlike traditional metrics and profit-based alternatives such as EMP, e-Profits evaluates models based on individualised cost–revenue dynamics, enabling closer alignment between predictive analytics and financial outcomes. Empirical results on two benchmark datasets show that e-Profits induces profit-relevant model re-ranking: models that are suboptimal under Accuracy or AUC often emerge as financially superior, particularly when evaluation focuses on high-value customer segments. Overall, the consistent divergence between full-population and top-segment profitability has direct implications for retention campaign design.
Beyond empirical performance, e-Profits is designed for practical deployment. e-Profits provides a more decision-relevant basis for model selection. Its plug-in formulation allows it to be applied to any churn classifier using predicted probabilities, without modifying the learning algorithm, and it integrates naturally into standard evaluation and model-selection workflows such as cross-validated tuning. The transparent, per-customer profit decomposition further facilitates communication between technical teams and business stakeholders. Overall, these results demonstrate that traditional predictive metrics are insufficient proxies for business value in churn prediction, and that profit-aware evaluation using e-Profits provides a more decision-relevant basis for model selection.
[…]
Acknowledgments This publication has emanated from research conducted with the support of Research Ireland under award no. GOIPG/2021/1354, and with the financial support of Research Ireland under grant no. 13/RC/2106_P2 at the ADAPT Research Centre at Technological University Dublin.
Declaration
@@
Ethics approval and consent to participate
Not applicable
Data availability
Materials availability
Not applicable
Code availability
All source code is available at: https://github.com/Awaismanzoor/eprofits
Author contribution
Awais Manzoor: Conceptualization, Data curation, Methodology and Experimental setup, Software, Formal analysis, Investigation, Validation, Writing – original draft, Writing – review & editing M. Atif Qureshi: Conceptualization, Data curation, Methodology and Experimental setup, Software, Formal analysis, Investigation, Validation, Writing – review & editing, Supervision Etain Kidney: Supervision, Writing – review & editing Luca Longo: Validation
Appendix A Toy Example: e-Profits Calculation
A.1 e-Profits Calculation Using TRR (conditional one-period retention)
To illustrate how e-Profits evaluates profitability at the customer level, we present a toy example of five customers with known churn labels $y_{i}$, predictions $\hat{y}_{i}$, and financial attributes. Following the targeting policy in Eq. 2, only customers predicted as churners ($\hat{y}_{i}=1$) are considered for intervention.
In this example, the tenure-based retention probability (TRR) is interpreted as a conditional one-period retention probability. Let $S(\cdot)$ denote the Kaplan-Meier survival function estimated from churn durations. Given a customer with current tenure $T_{i}$, the conditional retention over a horizon $\Delta$ is
$r_{i}^{\text{TRR}}(\Delta)=\Pr(T>T_{i}+\Delta\mid T>T_{i})=\frac{S(T_{i}+\Delta)}{S(T_{i})}.$
For simplicity, we use $\Delta=1$ (one period, e.g., one month). The retention values shown in Table 9 should be understood as $r_{i}^{\text{TRR}}(1)$ (i.e., one-period conditional retention), obtained by evaluating the fitted Kaplan-Meier curve.
Table 9: Example customer data for e-Profits illustration (TRR values are conditional one-period retention probabilities)
Customer $y_{i}$ (True) $\hat{y}_{i}$ (Pred) $R_{i}$ (Revenue) $r_{i}^{\text{TRR}}(1)$ (Cond. retention) Account Length $T_{i}$
A 1 1 100 0.80 120
B 0 1 80 0.90 50
C 1 0 120 0.75 180
D 0 0 90 0.85 90
E 1 1 150 0.70 200
Let us assume the profit margin $M=0.3$, the cost-per-offer ratio $\text{CPO}=0.1$, and contact cost parameters $c_{0}=5$, $c_{1}=0.3$ [^4] The contact cost is determined as follows: $c_{0}$ represents the minimum cost incurred regardless of CLV (e.g., email costs, phone contact). This minimum cost is compared to a CLV-based proportionate contact cost ($c_{1}$ times the offer cost), and the greater value is selected.. Under TRR, CLV is computed as:
$\text{CLV}_{i}^{\text{TRR}}=\frac{R_{i}\cdot M}{1-\min\!\left(r_{i}^{\text{TRR}}(1),0.995\right)}.$
The intervention costs for each predicted churner are:
$C_{\text{offer},i}=\text{CPO}\cdot\text{CLV}_{i},\quad C_{\text{contact},i}=\max(c_{0},\,c_{1}\cdot C_{\text{offer},i}).$
Customer A (TP):
$\text{CLV}_{A}=\frac{100\cdot 0.3}{1-0.80}=150,\quad C_{\text{offer}}=15,\quad C_{\text{contact}}=\max(5,4.5)=5$
$\Pi_{A}=150-15-5=\mathbf{130}$
Customer B (FP):
$\text{CLV}_{B}=\frac{80\cdot 0.3}{1-0.90}=240,\quad C_{\text{offer}}=24,\quad C_{\text{contact}}=\max(5,7.2)=7.2$
$\Pi_{B}=-24-7.2=\mathbf{-31.2}$
Customer C and D:
$\hat{y}_{C}=\hat{y}_{D}=0\Rightarrow\Pi_{C}=\Pi_{D}=0$
Customer E (TP):
$\text{CLV}_{E}=\frac{150\cdot 0.3}{1-0.70}=150,\quad C_{\text{offer}}=15,\quad C_{\text{contact}}=\max(5,4.5)=5$
$\Pi_{E}=150-15-5=\mathbf{130}$
Total e-Profits (TRR-based):
$\sum\Pi_{i}^{\text{TRR}}=130+(-31.2)+0+0+130=\mathbf{228.8}.$
A.2 e-Profits Calculation Using ARR (global one-period retention)
To contrast tenure-based retention with a population-level assumption, we repeat the toy example using an Average Retention Probability (ARR). To remain consistent with the conditional TRR definition, we define ARR as the mean of the conditional one-period retention probabilities:
$r^{\text{ARR}}(1)=\frac{1}{N}\sum_{i=1}^{N}r_{i}^{\text{TRR}}(1).$
Using the five customers:
$r^{\text{ARR}}(1)=\frac{0.80+0.90+0.75+0.85+0.70}{5}=\frac{4.00}{5}=0.80.$
ARR-based CLV applies this global retention uniformly:
$\text{CLV}_{i}^{\text{ARR}}=\frac{R_{i}\cdot M}{1-\min\!\left(r^{\text{ARR}}(1),0.995\right)}.$
With $r^{\text{ARR}}(1)=0.80$, this becomes:
$\text{CLV}_{i}^{\text{ARR}}=\frac{0.3\cdot R_{i}}{1-0.80}=\frac{0.3\cdot R_{i}}{0.2}=1.5\cdot R_{i}.$
Intervention costs are computed identically:
$C_{\text{offer},i}=\text{CPO}\cdot\text{CLV}_{i},\quad C_{\text{contact},i}=\max(c_{0},\,c_{1}\cdot C_{\text{offer},i}).$
Customer A (TP):
$\text{CLV}_{A}=1.5\times 100=150,\quad C_{\text{offer}}=15,\quad C_{\text{contact}}=\max(5,4.5)=5$
$\Pi_{A}=150-15-5=\mathbf{130}$
Customer B (FP):
$\text{CLV}_{B}=1.5\times 80=120,\quad C_{\text{offer}}=12,\quad C_{\text{contact}}=\max(5,3.6)=5$
$\Pi_{B}=-12-5=\mathbf{-17}$
Customer C and D:
$\hat{y}_{C}=\hat{y}_{D}=0\Rightarrow\Pi_{C}=\Pi_{D}=0$
Customer E (TP):
$\text{CLV}_{E}=1.5\times 150=225,\quad C_{\text{offer}}=22.5,\quad C_{\text{contact}}=\max(5,6.75)=6.75$
$\Pi_{E}=225-22.5-6.75=\mathbf{195.75}$
Total e-Profits (ARR-based):
$\sum\Pi_{i}^{\text{ARR}}=130+(-17)+0+0+195.75=\mathbf{308.75}.$
This toy example shows how e-Profits combines prediction correctness with customer-level economics. Customers A and E contribute positively because they are true positives, while Customer B produces a loss as a false positive due to unnecessary intervention costs. The difference between TRR and ARR arises because TRR uses customer-specific conditional retention $r_{i}^{\text{TRR}}(1)$ whereas ARR applies a single averaged retention probability $r^{\text{ARR}}(1)$ uniformly.